# SCorP: Statistics-Informed Dense Correspondence Prediction Directly from Unsegmented Medical Images

### v1

@@ 1 Introduction
Figure 1 a and b provide an overview of the requirements, limitations, and visual comparison of different SSM pipelines. Notably, our proposed method SCorP distinguishes itself by its minimal requirements (Figure 1.a), relying solely on surface representation in the form of meshes, point clouds, or binary volumes for training while avoiding adherence to the linearity assumption. Figure 1.b further provides a visual comparison of the training pipelines of various methods in contrast to SCorP.
Our main contributions are:
@@ 2 Related Work
Deep learning models such as DeepSSM and TL-DeepSSM [7, 6] provide alternatives to traditional SSM tools and are gaining traction. These models learn a functional mapping parameterized by a deep network that estimates surface correspondences from unsegmented images in a supervised manner. Several models have been proposed to enhance the performance of DeepSSM [7, 6]. These modifications include incorporating multi-scale and progressive learning modules (e.g., Progressive DeepSSM [5]), introducing anatomy localization modules for raw images (e.g., LocalizedSSM [28]), and introducing uncertainty quantification (e.g., Uncertain DeepSSM [1], VIB-DeepSSM [2], BVIB-DeepSSM [4]). Despite these advancements, these models still rely on optimized PDMs for training.
Other deep learning-based image-driven SSM methods have been introduced that leverage radial basis functions (RBF)-based representation to learn control points and normals for surface estimation [30]. However, these models face scalability challenges with large datasets and increased correspondences required to model complex anatomies.
@@ 4.1 Datasets
Left Atrium Dataset (LA): The dataset comprises 923 anonymized Late Gadolinium Enhancement (LGE) MRIs obtained from distinct patients and were manually segmented by cardiovascular medicine experts. They have a spatial resolution of $0.65\times 0.65\times 2.5mm^{3}$, with the endocardial wall serving as the boundary for the pulmonary veins. Following segmentation, the images were cropped around the region of interest and downsampled by a factor of 0.8 to effectively manage memory usage, resulting in input images of size $166\times 120\times 125$.
@@ 4.2 Models for Comparison
    DeepSSM [6] is a leading supervised model for predicting correspondence points from 3D image volumes. This method necessitates an optimized PDM for training, where each training instance consists of an image-correspondence pair. We utilize the correspondence-supervised version of DeepSSM that uses a fixed decoder initialized with PCA basis and mean shape and trained on mean squared error (MSE) loss between predicted and ground truth correspondences.
@@ 4.4 Experimental Setup
We use the medoid shape of each dataset with 1024 correspondences as the template for the implicit field decoder. The medoid shape is identified using the surface-to-surface distances of meshes. This template remains consistent across all three training phases. We employ the Adam optimizer with a fixed learning rate of $0.00001$ and continue training until convergence, determined through validation evaluation. Convergence is reached when the validation CD does not improve for 200 epochs. The models resulting from the epoch with the best validation CD are chosen for evaluation. Additionally, we set the weighting term of the vertex reconstruction term $\alpha$ to $0.001$ (eq. 1) for all experiments. During the mesh branch training, a random jitter with a standard deviation of $1\%$ of the maximum vertex size is added as data augmentation. The hyperparameters are identified via tuning for the validation set performance.
@@ 7 Conclusion
The proposed framework, SCorP, presents a novel approach to inferring correspondences directly from raw images without needing a pre-optimized shape model. By integrating prior shape information from surface representations ( meshes, point clouds, binary volumes), SCorP achieves superior performance compared to traditional and state-of-the-art deep learning methods with less supervision. The three-phase training strategy ensures effective integration of shape statistics-informed priors, guiding the image encoder to learn representative shape features for the correspondence prediction task. Furthermore, SCorP demonstrates robustness across varying training dataset sizes, highlighting its versatility and applicability in different scenarios. Overall, SCorP improves upon existing methods by streamlining the PDM generation process, which increases the feasibility of using shape models for research and applications in medical imaging, computer-aided diagnosis, and beyond.

### v2

@@ 1 Introduction
Figure 1 a and b provide an overview of the requirements, limitations, and visual comparison of different SSM pipelines. Notably, our proposed method SCorP distinguishes itself by its minimal requirements (Figure 1.a), relying solely on surface representation in the form of meshes, point clouds, or binary volumes for training while avoiding adherence to the linearity assumption. Figure 1.b further provides a visual comparison of the training pipelines of various methods in contrast to SCorP. Our main contributions are:
@@ 2 Related Work
Deep learning models such as DeepSSM and TL-DeepSSM [7, 6] provide alternatives to traditional SSM tools and are gaining traction. These models learn a functional mapping parameterized by a deep network that estimates surface correspondences from unsegmented images in a supervised manner. Several models have been proposed to enhance the performance of DeepSSM [7, 6]. These modifications include incorporating multi-scale and progressive learning modules (e.g., Progressive DeepSSM [5]), introducing anatomy localization modules for raw images (e.g., LocalizedSSM [28]), and introducing uncertainty quantification (e.g., Uncertain DeepSSM [1], VIB-DeepSSM [2], BVIB-DeepSSM [4]). Despite these advancements, these models still rely on optimized PDMs for training. Other deep learning-based image-driven SSM methods have been introduced that leverage radial basis functions (RBF)-based representation to learn control points and normals for surface estimation [30]. However, these models face scalability challenges with large datasets and increased correspondences required to model complex anatomies.
@@ 4.1 Datasets
Left Atrium Dataset (LA): The dataset comprises 923 anonymized Late Gadolinium Enhancement (LGE) MRIs obtained from distinct patients and were manually segmented by cardiovascular medicine experts. The images were manually segmented at the University of Utah Division of Cardiovascular Medicine, the endocardium wall was used to cut off pulmonary veins. They have a spatial resolution of $0.65\times 0.65\times 2.5mm^{3}$, with the endocardial wall serving as the boundary for the pulmonary veins. Following segmentation, the images were cropped around the region of interest and downsampled by a factor of 0.8 to effectively manage memory usage, resulting in input images of size $166\times 120\times 125$.
@@ 4.2 Models for Comparison
    DeepSSM [6] is a leading supervised model for predicting correspondence points from 3D image volumes. This method necessitates an optimized PDM for training, where each training instance consists of an image-correspondence pair. We utilize the correspondence supervised version of DeepSSM that uses a fixed decoder initialized with PCA basis and mean shape and trained on mean squared error (MSE) loss between predicted and ground truth correspondences.
@@ 4.3 Metrics
This section describes the metrics used to assess the performance of the quality of the shape models. Since SCorP does not use the ground truth PDM for training, we exclude the root mean square error metric, which is used in DeepSSM [7] and TL-DeepSSM [7, 6].
@@ 4.4 Experimental Setup
We use the medoid shape of each dataset with 1024 correspondences as the template for the implicit field decoder. The medoid shape is identified using the surface-to-surface distances of meshes. This template remains consistent across all three training phases. We employ the Adam optimizer with a fixed learning rate of $0.00001$ and continue training until convergence, determined through validation evaluation. Convergence is reached when the validation CD does not improve for 200 epochs. The models resulting from the epoch with the best validation CD are chosen for evaluation. Additionally, we set the weighting term of the vertex reconstruction term $\alpha$ to $0.001$ (eq. 1) for all experiments. During the mesh branch training, a random jitter with a standard deviation of $1\%$ of the maximum vertex size is added as data augmentation. The hyperparameters are identified via tuning for the validation set performance. The source code is available at https://github.com/iyerkrithika21/SCorP_MIUA2024.
@@ 7 Conclusion
The proposed framework, SCorP presents a novel approach to inferring correspondences directly from raw images without needing a pre-optimized shape model. By integrating prior shape information from surface representations (meshes, point clouds, binary volumes), SCorP achieves superior performance compared to traditional and state-of-the-art deep learning methods with less supervision. The three-phase training strategy ensures effective integration of shape statistics-informed priors, guiding the image encoder to learn representative shape features for the correspondence prediction task. Furthermore, SCorP demonstrates robustness across varying training dataset sizes, highlighting its versatility and applicability in different scenarios. Overall, SCorP improves upon existing methods by streamlining the PDM generation process, which increases the feasibility of using shape models for research and applications in medical imaging, computer-aided diagnosis, and beyond.
8 Acknowledgements
This work was supported by the National Institutes of Health under grant numbers NIBIB-U24EB029011, NIAMS-R01AR076120, and NHLBI-R01HL135568. We thank the University of Utah Division of Cardiovascular Medicine for providing left atrium MRI scans and segmentations from the Atrial Fibrillation projects and the ShapeWorks team.